Find a function v such that f(z) = u+iv is analytic [i.e. find the conjugate function of u ].

1
Expert's answer
2021-09-02T00:21:56-0400

For analyticity in a region "R" of the "z" -plane, "f" must satisfy  Cauchy-Riemann equations relations


"u_x=v_y, u_y=-v_x"

at every point in that region.

Let "f(z) = u + i v" be an analytic function.

Given: "u(x, y)"

Then


"v=\int(-u_y)dx+\int u_xdy"